OpenStudy (anonymous):

Solve the system: x^2+y^2=5 x+y=3

OpenStudy (anonymous):

x=1 y=2

OpenStudy (anonymous):

express y in terms of x from 2º equation: y=3-x substitute in the 1º: x^2+(3-x)^2=5 solve: x^2+9-6x+x^2=5 2x^2-6x+4=0 x^2-3x+2=0 (x-2)(x-1)=0 so two posible values for x x=1, and x=2 if x=1 y=2 if x=2 y=1